# Yang-Baxter Hochschild Cohomology

### v4

@@ 2. Braided algebras
(1) $(\mu\otimes\mathbb{1})({\mathbb{1}}\otimes R)(R\otimes{\mathbb{1}})$ $=$ $R\otimes(\mathbb{1}\otimes\mu),$
(2) $(\mathbb{1}\otimes\mu)(R\otimes{\mathbb{1}})({\mathbb{1}}\otimes R)$ $=$ $R\otimes(\mu\otimes{\mathbb{1}}),$
respectively. We call $(V,\mu,R)$ a YI algebra (resp. IY algebra). If $(V,\mu,R)$ satisfies both Equation (1) and Equation (2), we call it a braided algebra.
@@ 2.1. Braided algebras from tensorized multiple conjugation quandles
A multiple conjugation quandle (MCQ) $X$ is the disjoint union of groups $G_{\lambda}$, where $\lambda$ is an element of an index set $\Lambda$, with a binary operation $*:X\times X\to X$ satisfying the following axioms.
@@
Let $X=\sqcup_{\lambda\in\Lambda}G_{\lambda}$ be an MCQ with a quandle operation *. Let ${\mathbb{k}}[X]$ be the free ${\mathbb{k}}$-module generated by $X$. Define a multiplication $\mu:{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]\rightarrow{\mathbb{k}}[X]$ on generators by $\mu(x\otimes y):=xy$ if $x,y\in G_{\lambda}$ for some $\lambda\in\Lambda$, and $\mu(x\otimes y):=0$ otherwise, and extended linearly. Define $R:{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]\rightarrow{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]$ on generators $x,y\in X$ by $R(x\otimes y):=y\otimes(x*y)$ and extended linearly. Then ${\mathbb{k}}[X]$ is a braided algebra, as we now proceed to show. If two of $x,y,z\in X$ belong to distinct $G_{\lambda}$’s, then the triple product is 00, and if all belong to the same $G_{\lambda}$ then it is associative, hence $\mu$ is associative. Equations (1) and (2), respectively, follow from the conditions $(ab)*x=(a*x)(b*x)$ and $x*(ab)=(x*a)*b$ for any $x\in X$, $a,b\in G_{\lambda}$ on generators, and if $a,b$ belong to distinct $G_{\lambda}$ then the values are 00 for both sides of the equations, hence ${\mathbb{k}}[X]$ is a YI and IY algebra.
@@
$R((x,y)\otimes(u,v)):=(u,v)\otimes[(x,y)*(u,v)]=(u,v)\otimes(xu^{-1},yu^{-1}v)$
@@ 2.3. Braided algebras from Hopf algebras
We now describe an important class of braided algebras related to MCQs derived from Hopf algebras. First, recall [EZ] that given a Hopf algebra $H$, one has a Yang-Baxter operator $R_{H}$ associated to it defined through the left adjoint action of $H$ on itself, which is a generalization of the notion of conjugation in a group:
@@
We show that Equation (2) holds on simple tensors. A similar procedure can be performed for Equation (1). For the left hand side of Equation (2) we have
$(\mu\otimes\mathbb{1})(\mathbb{1}\otimes R_{H})(R_{H}\otimes\mathbb{1})(x\otimes y\otimes z)$ $=$ $(\mu\otimes\mathbb{1})(\mathbb{1}\otimes R_{H})(y^{(1)})\otimes S(y^{(2)})xy^{(3)}\otimes z$
$=$ $(\mu\otimes\mathbb{1})(y^{(1)})\otimes z^{(3)}\otimes S(z^{(2)})S(y^{(2)})xy^{(3)}z^{(3)}$
@@ 3.1. Hochschild cohomology
Let $(V,\mu)$ be an associative algebra, $\tilde{V}=V\otimes_{\mathbb{k}}{\mathbb{k}}[[\hbar]]/(\hbar^{2})\cong V\oplus\hbar V$, and $\psi\in Z^{2}_{\rm H}(V,V)$. Then, setting $\tilde{\mu}=\mu+\hbar\psi$, $(\tilde{V},\tilde{\mu})$ is an algebra if and only if the equations $\delta^{2}_{\rm H}(\psi)=0$ hold. This is seen by computing the associativity in $\tilde{V}$,
@@
$\tilde{\mu}({\mathbb{1}}\otimes\tilde{\mu})$ $=$ $(\mu+\hbar\psi)({\mathbb{1}}\otimes(\mu+\hbar\psi)$
@@ 3.2. Yang-Baxter cohomology up to dimension 3
Let $f$ be a YB $1$-cochain. Then, the lemma is proved by showing that $\delta^{2}_{\rm YB}(\phi)=0$ when $\phi=\delta^{1}_{\rm YB}(f)$. Diagrammatically, when $\delta^{1}_{\rm YB}(f)$ is substituted in $\phi$, the circles crossing representing $\phi$ are replaced by four terms with circles placed at the four edges adjacent to the crossing. For example, the first term of $\delta^{2}_{\rm YB}(\phi)$ is shown in Figure 6 left top, and the replacement representing the substitution is depicted in the right of the top row. Similar diagrams are shown in the middle row for the second term of $\delta^{2}_{\rm YB}(\phi)$. One sees that the terms labeled (B) and (C) in the top and the middle row cancel with opposite signs. This cancelation principle is explained by the fact that each edge is shared by two crossings. In Figure 6, the terms labeled (A) in the top row and (D) in the bottom row have the circles for $f$ placed at the end edge (top). They are canceled by opposite signs after applying the YBE (in diagrammatic form). Hence in both cases (the circle placed inside or end edges), two terms cancel with opposite signs, and we obtain $\delta^{2}_{\rm YB}(\phi)=0$. ∎
@@
This YB cohomology theory is different from those in [Eisermann]; the differentials in [Eisermann] differ from ours by compositions of YB operators. For example, the 2-cochain $c^{\prime}$ in [Eisermann] is written as $Rc$ for our 2-cochain $c$. We adopt our definition and diagrammatics along the line of [CCES-coalgebra, CCES-adjoint] for the purpose of unifying Yang-Baxter cohomology with algebra cohomology.
@@ 3.3. Yang-Baxter Hochschild cohomology up to dimension $3$
The second differential $\delta^{2}_{\rm YBH}$ is defined to be the direct sum of four terms $\delta^{2}_{\rm YB}\oplus\delta^{2}_{\rm YI}\oplus\delta^{2}_{\rm IY}\oplus\delta^{2}_{\rm H}$ where $\delta^{2}_{\rm YB}$ and $\delta^{2}_{\rm H}$ map in the first ($C^{3,3}_{\rm YBH}(V,V)={\rm Hom}(V^{\otimes 3},V^{\otimes 3})$) and last $(C^{3,1}_{\rm YBH}(V,V)={\rm Hom}(V^{\otimes 3},V)$) copies of $C^{3}_{\rm YBH}(V,V)$, respectively, while $\delta^{2}_{\rm YI}$ and $\delta^{2}_{\rm IY}$ map to the middle two factors $C^{3,2}_{\rm YI}(V,V)$ and $C^{3,2}_{\rm IY}(V,V)$, respectively. Each differential is defined as follows.
@@
Let $f$ be a YBH $1$-cochain. Setting $\psi=\delta^{1}_{\rm H}(f)$ and $\phi=\delta^{1}_{\rm YB}(f)$, we show that $\delta^{2}_{i}(\phi\oplus\psi)=0$ for $i={\rm YI},{\rm IY}$. The procedure showing that $\delta^{2}_{\rm YI}(\phi\oplus\psi)=0$ is depicted in Figures 8 and 9, and similar to the proof of Lemma 3.1. In the right column of Figure 8, the positive terms of $\delta^{2}_{\rm IY}(\phi\oplus\psi)$ depicted in Figure 7 are shown. Circled vertices represent $\psi$ (3-valent vertices) and $\phi$ (4-valent ones). Under the assumption $\psi=\delta^{1}_{\rm H}(f)$ and $\phi=\delta^{1}_{\rm YB}(f)$, each vertex is replaced by a linear combination of the diagrams with circles (representing $f$) on edges, according to the definition of $\delta^{1}(f)$ (Figure 2). Such linear combinations for each term in the left column of Figure 8 are depicted in the right column of the figure. Thus $\psi=\delta^{1}_{\rm H}(f)$ and $\phi=\delta^{1}_{\rm YB}(f)$ is represented by replacing the positive terms in Figure 7 by the right columns in Figure 8. The negative terms are similarly replaced by the right columns in Figure 9. With signs one sees that all terms cancel, indicating $\delta^{2}_{\rm IY}(\phi\oplus\psi)=0$. The case $\delta^{2}_{\rm YI}(\phi\oplus\psi)=0$ is similar. The other differentials for YBE and Hochschild are similar as well. ∎
@@ 3.4. Infinitesimal deformations of braided algebras
Let $(V,\mu,R)$ be a braided algebra, and let $\psi$ and $\phi$ denote an algebra $2$-cocycle and a YB $2$-cocycle, respectively. Then, setting $\tilde{\mu}=\mu+\hbar\psi$ and $\tilde{R}=R+\hbar\phi$, $(\tilde{V},\tilde{\mu},\tilde{R})$ is an infinitesimal braided algebra deformation of $(V,\mu,R)$ if and only if the equations $\delta^{2}_{\rm YBH}(\phi\oplus\psi)=0$ hold.
@@
Since $\psi$ and $\phi$ are algebra and YB $2$-cocycles, respectively, it follows that $\tilde{\mu}$ is associative, and that $\tilde{R}$ is a YB operator. Now, we need to show that $\tilde{\mu}$ and $\tilde{R}$ satisfy the defining axioms of YI and IY algebras if and only if they satisfy the equations given in the statement. Let us consider the equation $(\tilde{\mu}\otimes\mathbb{1})(\mathbb{1}\otimes\tilde{R})(\tilde{R}\otimes\mathbb{1})=\tilde{R}\otimes(\mathbb{1}\otimes\tilde{\mu})$ in $\tilde{V}=V\otimes{\mathbb{k}}[[\hbar]]/(\hbar^{2})$. We have for Equation (2)
@@
$\tilde{R}\otimes(\mathbb{1}\otimes\tilde{\mu})$ $=$ $R\otimes(\mathbb{1}\otimes\mu)+\hbar\phi\otimes(\mathbb{1}\otimes\mu)+\hbar R\otimes(\mathbb{1}\otimes\psi).$
Using the fact that $(V,\mu,R)$ is an IY algebra, we find that $(\tilde{\mu}\otimes\mathbb{1})(\mathbb{1}\otimes\tilde{R})(\tilde{R}\otimes\mathbb{1})=\tilde{R}\otimes(\mathbb{1}\otimes\tilde{\mu})$ holds true if and only if $\delta^{2}_{\rm IY}(\phi\oplus\psi)=0$ is satisfied. One can then proceed analogously to show that the equation $(\mathbb{1}\otimes\tilde{\mu})(\tilde{R}\otimes\mathbb{1})(\mathbb{1}\otimes\tilde{R})=\tilde{R}\otimes(\tilde{\mu}\otimes\mathbb{1})$ if and only if $\delta^{2}_{\rm YI}(\phi\oplus\psi)=0$ holds, up to higher order terms in $\hbar$. ∎
@@
The two terms are seen to be equal since $\delta^{1}_{\rm YB}f=\phi$. A similar inspection of the homomorphism condition for associative algebras shows also that $\tilde{f}\tilde{\mu}=\mu(\tilde{f}\otimes\tilde{f})$. A similar procedure shows that the inverse of $\tilde{f}$ is a braided algebra homomorphism, completing the first part of the proof.
@@ 4. Braided multiplication and deformation cohomology
In this and the next sections, we discuss nontriviality of the second Yang-Baxter Hochschild cohomology groups. The third YBH cohomology groups are defined in Section 6 and independent from these two sections. The reader who would like to proceed to the third cohomology groups can go directly to Section 6.
@@
For $(\phi,\psi)\in C^{2}_{\rm YBH}(V,V)$ of $(V,\mu,R)$, define 2-cochains $(\phi_{R},\psi_{R})$ of $(V,\mu R,R)$ by $\psi_{R}:=\mu\phi+\psi R$ and $\phi_{R}:=\phi$, $(\psi_{R},\phi_{R})\in C^{2}_{\rm YBH}(V_{R},V_{R})$, as well as $f_{R}\in C^{1}_{\rm YBH}(V_{R},V_{R})$ by $f_{R}:=f$ for $f\in C^{1}_{\rm YBH}(V,V)$.
First we show that $(\phi_{R},\psi_{R})\in Z^{2}_{\rm YBH}(V_{R},V_{R})$. Each of associativity, Equation (1), and Equation (2) for $V_{R}$ follow from those of $V$. In the infinitesimal deformation $\tilde{V}$ of $V$, $(\mu+\hbar\psi,R+\hbar\phi)$ satisfies these relations, from the 2-cocycle relations of $(\phi,\psi)$. In order to show that $(\phi_{R},\psi_{R})$ satisfies the 2-cocycle conditions, it is sufficient to show that the infinitesimal deformation $\tilde{V}_{R}$ of $V_{R}$, $(\mu_{R}+\hbar\psi_{R},R+\hbar\phi_{R})$ satisfy the associativity, Equation (1) and Equation (2). We note that
@@
and $R+\hbar\phi_{R}=R+\hbar\phi$. Therefore the fact that the infinitesimal deformation $(\mu_{R}+\hbar\psi_{R},R+\psi_{R})$ giving a braided algebra modulo (2)superscriptPlanck-constant-over-$(\hbar^{2})$ follows from the assumption that $(\mu+\hbar\psi,R+\hbar\phi)$ giving the infinitesimal deformation that is a braided algebra. Hence the claim follows.
@@
as desired. It is also clear that $\iota_{R}$ is a $\mathbb{k}$-module homomorphism. To show injectivity, assume that $(\phi_{R},\psi_{R})$ is a coboundary, that is, $\psi_{R}=\delta^{1}_{\rm H}(f)$ and $\phi_{R}=\delta^{1}_{\rm YB}(f)$. Since $\phi_{R}=\phi$, we have $\phi=\delta^{1}_{\rm YB}(f)$. From $\psi_{R}=\delta^{1}_{\rm H}(f)$ we obtain $\phi\mu+\psi R=\mu R(f\otimes{\mathbb{1}})+\mu R({\mathbb{1}}\otimes f)-f\mu R$. Substituting $\phi=\delta^{1}_{\rm YB}(f)$ in the left hand side, we obtain

### v5

@@ 2. Braided algebras
(1) $(\mathbb{1}\otimes\mu)(R\otimes{\mathbb{1}})({\mathbb{1}}\otimes R)$ $=$ $R(\mu\otimes{\mathbb{1}}),$
(2) $(\mu\otimes\mathbb{1})({\mathbb{1}}\otimes R)(R\otimes{\mathbb{1}})$ $=$ $R(\mathbb{1}\otimes\mu),$
respectively. We call $(V,\mu,R)$ a YI algebra (resp. IY algebra), the naming defined in [JSC]. If $(V,\mu,R)$ satisfies both Equation (1) and Equation (2), we call it a braided algebra.
@@ 2.1. Braided algebras from tensorized multiple conjugation quandles
A multiple conjugation quandle (MCQ) $X$ is the disjoint union of groups $G_{\lambda}$, where $\lambda$ is an element of a finite index set $\Lambda$, with a binary operation $*:X\times X\to X$ satisfying the following axioms.
@@
Let $X=\sqcup_{\lambda\in\Lambda}G_{\lambda}$ be an MCQ with a quandle operation *. Let ${\mathbb{k}}[X]$ be the free ${\mathbb{k}}$-module generated by $X$. Define a multiplication $\mu:{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]\rightarrow{\mathbb{k}}[X]$ on generators by $\mu(x\otimes y):=xy$ if $x,y\in G_{\lambda}$ for some $\lambda\in\Lambda$, and $\mu(x\otimes y):=0$ otherwise, and extended linearly. Define $R:{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]\rightarrow{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]$ on generators $x,y\in X$ by $R(x\otimes y):=y\otimes(x*y)$ and extended linearly. Then ${\mathbb{k}}[X]$ is a braided algebra, as we now proceed to show. From the construction, it is seen that $R$ is a YBO. If two of $x,y,z\in X$ belong to distinct $G_{\lambda}$’s, then the triple product is 00, and if all belong to the same $G_{\lambda}$ then it is associative, hence $\mu$ is associative. Equations (1) and (2), respectively, follow from the conditions $(ab)*x=(a*x)(b*x)$ and $x*(ab)=(x*a)*b$ for any $x\in X$, $a,b\in G_{\lambda}$ on generators, and if $a,b$ belong to distinct $G_{\lambda}$ then the values are 00 for both sides of the equations, hence ${\mathbb{k}}[X]$ is a YI and IY algebra.
@@
$R((x,y)\otimes(u,v)):=(u,v)\otimes[(x,y)*(u,v)]=(u,v)\otimes(xu^{-1}v,yu^{-1}v)$
@@ 2.3. Braided algebras from Hopf algebras
We now describe an important class of braided algebras related to MCQs derived from Hopf algebras. First, recall [EZ] that given a Hopf algebra $H$, one has a Yang-Baxter operator $R_{H}$ associated to it defined through the right adjoint action of $H$ on itself, which is a generalization of the notion of conjugation in a group:
@@
We show that Equation (2) holds. A similar procedure can be performed for Equation (1). For the left hand side of Equation (2) we have
$(\mu\otimes\mathbb{1})(\mathbb{1}\otimes R_{H})(R_{H}\otimes\mathbb{1})(x\otimes y\otimes z)$ $=$ $(\mu\otimes\mathbb{1})(\mathbb{1}\otimes R_{H})(y^{(1)}\otimes S(y^{(2)})xy^{(3)}\otimes z)$
$=$ $(\mu\otimes\mathbb{1})(y^{(1)}\otimes z^{(1)}\otimes S(z^{(2)})S(y^{(2)})xy^{(3)}z^{(3)})$
@@ 3.1. Hochschild cohomology
Let $(V,\mu)$ be an associative algebra, $\tilde{V}=V\otimes_{\mathbb{k}}{\mathbb{k}}[[\hbar]]/(\hbar^{2})\cong V\oplus\hbar V$, and $\psi\in Z^{2}_{\rm H}(V,V)$. Then, setting $\tilde{\mu}=\mu+\hbar\psi$, $(\tilde{V},\tilde{\mu})$ is an associative algebra if and only if the equations $\delta^{2}_{\rm H}(\psi)=0$ hold. This is seen by computing the associativity in $\tilde{V}$,
@@
$\tilde{\mu}({\mathbb{1}}\otimes\tilde{\mu})$ $=$ $(\mu+\hbar\psi)({\mathbb{1}}\otimes(\mu+\hbar\psi))$
@@ 3.2. Yang-Baxter cohomology up to dimension 3
Let $f$ be a YB $1$-cochain. Then, the lemma is proved by showing that $\delta^{2}_{\rm YB}(\phi)=0$ when $\phi=\delta^{1}_{\rm YB}(f)$. Diagrammatically, when $\delta^{1}_{\rm YB}(f)$ is substituted in $\phi$, the circles crossing representing $\phi$ are replaced by four terms with circles placed at the four edges adjacent to the crossing. For example, the first term of $\delta^{2}_{\rm YB}(\phi)$ is shown in Figure 6 left top, and the replacement representing the substitution is depicted in the right of the top row. Similar diagrams are shown in the middle row for the second term of $\delta^{2}_{\rm YB}(\phi)$, and the bottom row in Figure 6 comes from the fifth summand in Figure 5. One sees that the terms labeled (B) and (C) in the top and the middle row cancel with opposite signs. This cancelation principle is explained by the fact that each edge is shared by two crossings. In Figure 6, the terms labeled (A) in the top row and (D) in the bottom row have the circles for $f$ placed at the end edge (top). They are canceled by opposite signs after applying the YBE (in diagrammatic form). Hence in both cases (the circle placed inside or end edges), two terms cancel with opposite signs, and we obtain $\delta^{2}_{\rm YB}(\phi)=0$. ∎
@@
This YB cohomology theory is different from those in [Eisermann]; the differentials in [Eisermann] differ from ours by compositions of YB operators. For example, the 2-cochain $c^{\prime}$ in [Eisermann] is written as $Rc$ for a 2-cochain $c$ in the cohomology considered in this article. We adopt our definition and diagrammatics along the line of [CCES-coalgebra, CCES-adjoint] for the purpose of unifying Yang-Baxter cohomology with algebra cohomology.
@@ 3.3. Yang-Baxter Hochschild cohomology up to dimension $3$
The second differential $\delta^{2}_{\rm YBH}$ is defined to be the direct sum of four terms $\delta^{2}_{\rm YB}\oplus\delta^{2}_{\rm YI}\oplus\delta^{2}_{\rm IY}\oplus\delta^{2}_{\rm H}$ where $\delta^{2}_{\rm YB}$ and $\delta^{2}_{\rm H}$ map in the first ($C^{3,3}_{\rm YBH}(V,V)={\rm Hom}(V^{\otimes 3},V^{\otimes 3})$) and last $(C^{3,1}_{\rm YBH}(V,V)={\rm Hom}(V^{\otimes 3},V)$) direct summands of $C^{3}_{\rm YBH}(V,V)$, respectively, while $\delta^{2}_{\rm YI}$ and $\delta^{2}_{\rm IY}$ map to the middle two direct summands $C^{3,2}_{\rm YI}(V,V)$ and $C^{3,2}_{\rm IY}(V,V)$, respectively. Each differential is defined as follows.
@@
Let $f$ be a YBH $1$-cochain. Setting $\psi=\delta^{1}_{\rm H}(f)$ and $\phi=\delta^{1}_{\rm YB}(f)$, we show that $\delta^{2}_{i}(\phi\oplus\psi)=0$ for $i={\rm YI},{\rm IY}$. The procedure showing that $\delta^{2}_{\rm IY}(\phi\oplus\psi)=0$ is depicted in Figures 8 and 9, and similar to the proof of Lemma 3.1. In the right column of Figure 8, the positive terms of $\delta^{2}_{\rm IY}(\phi\oplus\psi)$ depicted in Figure 7 are shown. Circled vertices represent $\psi$ (3-valent vertices) and $\phi$ (4-valent ones). Under the assumption $\psi=\delta^{1}_{\rm H}(f)$ and $\phi=\delta^{1}_{\rm YB}(f)$, each vertex is replaced by a linear combination of the diagrams with circles (representing $f$) on edges, according to the definition of $\delta^{1}(f)$ (Figure 2). Such linear combinations for each term in the left column of Figure 8 are depicted in the right column of the figure. Thus $\psi=\delta^{1}_{\rm H}(f)$ and $\phi=\delta^{1}_{\rm YB}(f)$ is represented by replacing the positive terms in Figure 7 by the right columns in Figure 8. The negative terms are similarly replaced by the right columns in Figure 9. With signs one sees that all terms cancel, indicating $\delta^{2}_{\rm IY}(\phi\oplus\psi)=0$. The case $\delta^{2}_{\rm YI}(\phi\oplus\psi)=0$ is similar. The other differentials for YBE and Hochschild are similar as well. ∎
@@ 3.4. Infinitesimal deformations of braided algebras
Let $(V,\mu,R)$ be a braided algebra, and let $\psi$ and $\phi$ denote an algebra $2$-cocycle and a YB $2$-cocycle, respectively. Then, setting $\tilde{\mu}=\mu+\hbar\psi$ and $\tilde{R}=R+\hbar\phi$, $(\tilde{V},\tilde{\mu},\tilde{R})$ is an infinitesimal braided algebra deformation of $(V,\mu,R)$ if and only if $\delta^{2}_{\rm YBH}(\phi\oplus\psi)=0$ hold.
@@
Since $\psi$ and $\phi$ are algebra and YB $2$-cocycles, respectively, it follows that $\tilde{\mu}$ is associative, and that $\tilde{R}$ is a YB operator. Now, we need to show that $\tilde{\mu}$ and $\tilde{R}$ satisfy the defining axioms of YI and IY algebras if and only if they satisfy $\delta^{2}_{\rm YBH}(\phi\oplus\psi)=0$, as given in the statement. Let us consider the equation $(\tilde{\mu}\otimes\mathbb{1})(\mathbb{1}\otimes\tilde{R})(\tilde{R}\otimes\mathbb{1})=\tilde{R}(\mathbb{1}\otimes\tilde{\mu})$ in $\tilde{V}=V\otimes{\mathbb{k}}[[\hbar]]/(\hbar^{2})$. We have for Equation (2)
@@
$\tilde{R}(\mathbb{1}\otimes\tilde{\mu})$ $=$ $R(\mathbb{1}\otimes\mu)+\hbar\phi(\mathbb{1}\otimes\mu)+\hbar R(\mathbb{1}\otimes\psi).$
Using the fact that $(V,\mu,R)$ is an IY algebra, we find that $(\tilde{\mu}\otimes\mathbb{1})(\mathbb{1}\otimes\tilde{R})(\tilde{R}\otimes\mathbb{1})=\tilde{R}(\mathbb{1}\otimes\tilde{\mu})$ holds true if and only if $\delta^{2}_{\rm IY}(\phi\oplus\psi)=0$ is satisfied. One can then proceed analogously to show that the equation $(\mathbb{1}\otimes\tilde{\mu})(\tilde{R}\otimes\mathbb{1})(\mathbb{1}\otimes\tilde{R})=\tilde{R}(\tilde{\mu}\otimes\mathbb{1})$ if and only if $\delta^{2}_{\rm YI}(\phi\oplus\psi)=0$ holds, up to higher order terms in $\hbar$. ∎
@@
The two terms are seen to be equal since $\delta^{1}_{\rm YB}f=\phi$. A similar inspection of the homomorphism condition for associative algebras shows also that $\tilde{f}\tilde{\mu}=\mu(\tilde{f}\otimes\tilde{f})$. Analogously, one shows that the inverse of $\tilde{f}$ is a braided algebra homomorphism, completing the first part of the proof.
@@ 4. Braided multiplication and deformation cohomology
In this and the next section, we discuss nontriviality of the second Yang-Baxter Hochschild cohomology groups. The third YBH cohomology groups are defined in Section 6 and independent from these two sections. The reader who would like to proceed to the third cohomology groups can go directly to Section 6.
@@
For $(\phi,\psi)\in C^{2}_{\rm YBH}(V,V)$ of $(V,\mu,R)$, define 2-cochains $(\phi_{R},\psi_{R})$ of $(V,\mu R,R)$ by $\psi_{R}:=\mu\phi+\psi R$ and $\phi_{R}:=\phi$, $(\psi_{R},\phi_{R})\in C^{2}_{\rm YBH}(V_{R},V_{R})$, as well as $f_{R}\in C^{1}_{\rm YBH}(V_{R},V_{R})$ by $f_{R}:=f$ for $f\in C^{1}_{\rm YBH}(V,V)$. The map $\iota_{R}$ is defined by these assignments.
First we show that $(\phi_{R},\psi_{R})\in Z^{2}_{\rm YBH}(V_{R},V_{R})$. Each of associativity, Equation (1), and Equation (2) for $V_{R}$ follow from those of $V$, as shown in Figure (10). Analogously, in the infinitesimal deformation $\tilde{V}$ of $V$, $(\mu+\hbar\psi,R+\hbar\phi)$ satisfies these relations, from the 2-cocycle relations of $(\phi,\psi)$. In order to show that $(\phi_{R},\psi_{R})$ satisfies the 2-cocycle conditions, it is sufficient to show that the infinitesimal deformation $\tilde{V}_{R}$ of $V_{R}$, $(\mu_{R}+\hbar\psi_{R},R+\hbar\phi_{R})$ satisfies the associativity condition, Equation (1) and Equation (2). We note that
@@
and $R+\hbar\phi_{R}=R+\hbar\phi$. Therefore the fact that the infinitesimal deformation $(\mu_{R}+\hbar\psi_{R},R+\hbar\psi_{R})$ giving a braided algebra modulo (2)superscriptPlanck-constant-over-$(\hbar^{2})$ follows from the assumption that $(\mu+\hbar\psi,R+\hbar\phi)$ giving the infinitesimal deformation that is a braided algebra. Hence the claim follows.
@@
as desired. It is also clear that $\iota_{R}$ is a $\mathbb{k}$-module homomorphism. To show injectivity, assume that $(\phi_{R},\psi_{R})$ is a coboundary, that is, $\psi_{R}=\delta^{1}_{\rm H}(f)$ and $\phi_{R}=\delta^{1}_{\rm YB}(f)$. Since $\phi_{R}=\phi$, we have $\phi=\delta^{1}_{\rm YB}(f)$. From $\psi_{R}=\delta^{1}_{\rm H}(f)$ we obtain $\mu\phi+\psi R=\mu R(f\otimes{\mathbb{1}})+\mu R({\mathbb{1}}\otimes f)-f\mu R$. Substituting $\phi=\delta^{1}_{\rm YB}(f)$ in the left hand side, we obtain